OpenStudy (anonymous):

The lateral area of a cone is 480 pie cm^2.The radius is 40 cm.Find the slant height to the nearest tenth. 12cm 11.5cm 7.7cm 6.7cm SA = πr^2 + πrl

OpenStudy (anonymous):

SA= 3.14*40^2 + 3.14*40*480 1602.1 +125.6 = 1727.7 doesn't match any of the choices

OpenStudy (anonymous):

I got 12.

OpenStudy (anonymous):

The reason why it doesn't match up is since you have the formula of the normal surface area. πr² is the base, and lateral area does NOT include the base. Get rid of that and then solve for L. 40πl = 480, but still this doesn't work. since l ≈ 3.81

OpenStudy (anonymous):

LA = pi(r)l

OpenStudy (anonymous):

so 480 =rl, or 480 = 40l. Then divide 480/40 so that l = 12